If y = x^2 - 2x - 8 has two intersection points A and B with the x-axis, and its vertex is P. Find the area of ΔPAB.
Answer #1

sol. Given parabole y=%-20-8 y = (1-1),–8 =61-2²-9 (1-1)² = 4+9 (a- h ua (4-k) a represents a open up parabole whose verlex a